OpenStudy (anonymous):

i need help solving quadratic equations by factors. the equation is 5x^2=18x+35

OpenStudy (anonymous):

set equal to zero: 5x^2-18x-35=0 then find two numbers that multiply to get -175 and add to get -18, which are -25 and 7. use them to split the middle term: 5x^2-25x+7x-35=0 then use grouping: 5x(x-5)+7(x-5)=0 so (5x+7)(x-5)=0 setting each equal to zero, x=-7/5 or x=5

OpenStudy (anonymous):

do you need any more detail?

OpenStudy (anonymous):

no i dnt thank you for your help